OpenStudy (anonymous):

solve this equation x+2y+2z=6, -2x+4y+6z=-3, x+3y+z=6

OpenStudy (anonymous):

x + 2y + 2z = 6 -->(-1) x + 3y + z = 6 ------------- -x - 2y - 2z = -6 (result of multiplying by -1) x + 3y + z = 6 -------------add y - z = 0 x + 2y + 2z = 6 -->(2) -2x + 4y + 6z = -3 ------------- 2x + 4y + 4z = 12 (result of multiplying by 2) -2x + 4y + 6z = -3 -------------add 8y + 10z = 9 y - z = 0 --->(-8) 8y + 10z = 9 ------------- -8y + 8z = 0 (result of multiplying by -8) 8y + 10z = 9 ------------add 18z = 9 z = 1/2 y - z = 0 y - 1/2 = 0 y = 1/2 x + 2y + 2z = 6 x + 2(1/2) + 2(1/2) = 6 x + 1 + 1 = 6 x + 2 = 6 x = 6 - 2 x = 4 check.. x + 3y + z = 6 4 + 3(1/2) + 1/2 = 6 4 + 3/2 + 1/2 = 6 4 + 4/2 = 6 4 + 2 = 6 6 = 6 (correct) x = 4, y = 1/2 and z = 1/2